OpenStudy (anonymous):

I need help with this step by step Evaluate the indefinite integral. \int \frac{(\ln(x))^4}{x} dx by substitution

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ \ln(x)^4 }{ x }dx\]

OpenStudy (anonymous):

try \(u=ln(x), du =\frac{dx}{x}\) and you should get it in one step

OpenStudy (anonymous):

May I help?

OpenStudy (anonymous):

visualize it as \[\int\frac{ \ln(x)^4 }{ x }dx=\int \ln^4(x)\frac{dx}{x}\]

OpenStudy (anonymous):

pretend I'm a complete moron

OpenStudy (anonymous):

(x)^4 ?

OpenStudy (anonymous):

are you familiar with "u - substitution"

OpenStudy (anonymous):

\(\ln x^4\) Why \(\ln ^4 x\) @satellite73

OpenStudy (anonymous):

sorry, it's the whole thing to the 4th

OpenStudy (anonymous):

it is the chain rule backwards

OpenStudy (anonymous):

I understand that you pick one component to make u and that the derivative of that needs to be in the function it's what you do with it that I can't figure out

OpenStudy (anonymous):

first off, lets get the answer and see why it works. what do you get when you take the derivative of \(\ln^3(x)\) ? by the chain rule you get \(3\ln(x)\times \frac{1}{x}=\frac{3\ln^2(x)}{x}\)

OpenStudy (anonymous):

Ok! \((\ln x)^4\) is what we have

OpenStudy (anonymous):

this tells you \[\int\frac{3\ln^2(x)}{x}dx=\ln^3(x)\]

OpenStudy (anonymous):

yes @SheldonEinstein

OpenStudy (anonymous):

I would not interupt and let @satellite73 answer.

OpenStudy (anonymous):

so a u-sub is a way to undo the chain rule you have a composite function \(\ln^4(x)=(\ln(x))^4\) multiplied by the derivative of the "inside function" \(\frac{1}{x}\)

OpenStudy (anonymous):

if you make the substitution \(u=\ln(x)\) then the derivative is \(\frac{1}{x}\) so the gimmick is to say \[u=\ln(x), du=\frac{1}{x}dx\]

OpenStudy (anonymous):

then rewrite with only \(u\) terms and get \[\int u^4du\]

OpenStudy (anonymous):

Ok I'm with you so far

OpenStudy (anonymous):

now the anti derivative is relatively easy it is \(\frac{u^5}{5}\)

OpenStudy (anonymous):

putting the \(u=\ln(x)\) back we get \[\frac{\ln^5(x)}{5}\]

OpenStudy (anonymous):

why over 5? that's where I got lost

OpenStudy (anonymous):

check that this works by differentiation, and you will see that this is correct

OpenStudy (anonymous):

oh, over 5 because the derivative of \(u^5\) is \(5u^4\) but you want \(u^4\) and not \(5u^4\) so you have to divide by \(5\) to get it

OpenStudy (anonymous):

nevermind

OpenStudy (anonymous):

\[\int x^n dx =\frac{x^{n+1}}{n+1}\]

OpenStudy (anonymous):

I'm used to doing (1/5)u^5

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

check that the derivative of \[\frac{1}{5}\ln^5(x)\] is \[\frac{\ln^4(x)}{x}\] and you will see why the "u - substitution" works

OpenStudy (anonymous):

I did like this : \[\int \cfrac{ [\ln x]^4 dx }{x} \] Put \(\cfrac{1}{5} [\ln x]^5 = u \) \(\int du = u + c\) = \(\cfrac{1}{5} [\ln x]^5 + C\) \(\implies \cfrac{[\ln x]^4 dx}{x} = du\) Is this also correct @satellite73 ?

OpenStudy (anonymous):

it looks like you knew the answer at the start and then just said it look at this line \[\int du = u + c\] this says the integral is \(u\) which you knew at the beginning, probably by doing the u - sub in your head in other words, you knew what the answer was, but the u - sub is the way to get the answer

OpenStudy (anonymous):

both of those really help! I'll try the next one and see what happens

OpenStudy (anonymous):

the entire problem is to figure out the \[\frac{1}{5}\ln^5(x)\] part

OpenStudy (anonymous):

Ok! I got it. @heradog not me, @satellite73 was the one who helped. Best of luck for your journey and btw \(\mathbb{ I} \space \mathbb{LOVE} \space \mathbb{INTEGRALS} \)

OpenStudy (anonymous):

good i'm stuck again